What is the average acceleration of the particle within the time interval 1 second to 3 seconds?


1
Expert's answer
2020-12-21T11:00:44-0500

By definition, the average acceleration is


"a = \dfrac{\Delta v}{\Delta t}"

where "\Delta v" is the change of the particle's velocity during the time intarval "\Delta t". According to the ploblem statement, time interval is "\Delta t = 3s-1s = 2s". Thus, the acceleration is:


"a = \dfrac{\Delta v}{2}\space (m/s^2)"

for any velocity change "\Delta v\space (m/s)".


Answer. "a = \dfrac{\Delta v}{2}\space (m/s^2)" for any velocity change "\Delta v\space (m/s)"